# How Does Critical Batch Size Scale in Pre-training?

### v1

@@ 4.1 Fixed Data Size and Scaling Up Model Size
Various previous works have established infinite width limits of neural networks (Yang and Hu, 2021; Bordelon and Pehlevan, 2022). For initializations and architectures that obey these limits, we can theoretically claim, that for a fixed training time and batch size, the performance of the neural networks asymptotes with increasing width. The formal statement is provided below:
@@
For SGD with a given batch size $B$ (or for gradient descent, i.e., $B\rightarrow\infty$), training iterations $t$, an error tolerance $\epsilon>0$, fixed learning rate schedule and data ordering, for any network and initialization satisfying Master Theorem (Theorem G.4) in Yang and Hu (2021), there exists a width $w$ such that for any two networks $M_{1},M_{2}$ having widths $w_{1},w_{2}>w$, $|\mathcal{R}(M_{1},t)-\mathcal{R}(M_{2},t)|\leq\epsilon$, where $\mathcal{R}(M,t)$ denotes the loss of network $N$ at time $t$.
@@ 4.2 Fixed Model Size and Scaling up Data Size
We now turn to study the impact of the batch size in mini-batch SGD for a well-specified Gaussian linear regression problem. Let $(\mathbf{x},y)$ be a pair of covariates and responses from a population distribution. Let the population risk and the population distribution be
@@
where $\gamma>0$ is a constant learning rate, $B$ is the batch size, $n:=D/B$ is the number of steps, $\mathbf{w}_{0}=0$ is the initialization (without loss of generality), and the output is the average of the iterates, $\bar{\mathbf{w}}:=\frac{1}{n}\sum_{t=0}^{n-1}\mathbf{w}_{t}$. The following theorem characterizes the CBS of mini-batch SGD that minimizes the number of steps ($n$) without sacrificing the rate of the attained excess risk.
@@
We write $f(D)\lesssim g(D)$ if there is a positive constant $c$ such that $f(D)\leq cg(D)$ for every $D\geq 1$. We write $f(D)\eqsim g(D)$ if $f(D)\lesssim g(D)\lesssim f(D)$.
Assume the population distribution satisfies $\sigma^{2}\eqsim 1$ and the following capacity and source conditions:
@@
where $(\lambda_{i},\mathbf{v}_{i})_{i>0}$ are the eigenvalues and the corresponding eigenvectors of $\mathbf{H}$, and the expectation is over a prior of $\mathbf{w}^{*}$. Then for $b>a$, the CBS is $B^{*}\eqsim D^{1-a/\min\{b,2a+1\}}$, which allows the mini-batch SGD output $\bar{\mathbf{w}}$ to attain the optimal rate of the expected excess risk (as data size $D$ grows) with the smallest number of steps $n$.
Proof of Theorem 2.
It follows from Theorems 3 and 1 in Appendix G. ∎
The capacity and source conditions are from the nonparametric linear regression literature (Caponnetto and De Vito, 2007) and are recently used to study scaling laws theory (Bordelon et al., 2024c; Lin et al., 2024; Paquette et al., 2024). In the regime considered in Theorem 2, the variance error tends to dominate the bias error, and the optimal choices of batch size and learning rate balance these two errors. While one can use $B^{*}=1$ and a small $\gamma^{*}$ (see Corollary 1 in Appendix G) to achieve the best excess risk rate, this leads to a suboptimal sequential runtime ($n=D/B$). In this case, the CBS is $B^{*}\eqsim D^{1-a/\min\{b,2a+1\}}$, which achieves the optimal excess risk rate while minimizing the sequential runtime.
@@ 5 Related Work
Scaling laws. Scaling laws describe the parametric relationships among key factors involved in training neural networks: model size $N$, dataset size $D$, training $\operatorname{cost}C$, and final training loss $\mathcal{R}$. These laws enable the prediction of training loss $\mathcal{R}$ based on available resources, making it possible to optimize resource allocation for efficient model training. For example, Hestness et al. (2017) found $\mathcal{R}\propto D^{-\alpha}$, with $\alpha\in[0.07,0.35]$. Of the factors they varied, only task can change the exponent $\alpha$. Changing the architecture optimizers, regularizers, and loss functions, would only change the proportionality factor, not the exponent; Henighan et al. (2020) studied statistical relations between $N,D,C,\mathcal{R}$, over a wide range of values and found similar scaling laws, over the range of $N\in\left[10^{3},10^{9}\right],C\in\left[10^{12},10^{21}\right]$, and over multiple modalities (text, video, image, text to image, etc.). (Kaplan et al., 2020) states that $N$ should be scaled faster than $D$. However, Chinchilla scaling (Hoffmann et al., 2022b) found that models are under-trained, and then suggests that when given an increased budget (in FLOPs), to achieve compute-optimal, the number of model parameters ($N$) and the number of tokens for training the model ($D$) should scale in approximately equal proportions. Recent efforts (Pearce and Song, 2024; Besiroglu et al., 2024; Porian et al., 2024) have been made in reproducing the scaling laws from (Hoffmann et al., 2022b) and the (Kaplan et al., 2020). Different from our focus on measuring the efficiency notion of CBS, most of them focus on deriving optimal hyper-parameters (Bi et al., 2024; Porian et al., 2024) including learning rate and batch size from small-scale training given a fixed compute budget FLOPS $\approx 6ND$ without decoupling the effects of model size and data size.
@@
Optimization and critical batch size. Previous studies have shown that increasing batch sizes can be offset by a proportional adjustment to the learning rate in small-scale regimes (McCandlish et al., 2018; Zhang et al., 2019; Kaplan et al., 2020). McCandlish et al. (2018) introduced the gradient noise scale, a measure that captures the variation in gradients across different training examples, which helps predict the critical batch size (CBS). Their findings also suggest that small-batch training is more compute-efficient, while large-batch training requires fewer optimizer steps. Momentum-based methods extend scaling to larger batch sizes but converge to the performance of standard SGD at smaller batch sizes (Shallue et al., 2019). Additionally, Zhang et al. (2019) analyzed the impact of curvature on CBS using a noisy quadratic model, demonstrating that preconditioning techniques can increase the CBS. Golmant et al. (2018) showed that the size of the dataset plays a smaller role in determining training efficiency compared to factors like model architecture and data complexity. In contrast, Hilton et al. (2022) examined how performance can be maintained at smaller batch sizes. Meanwhile, Smith et al. (2017); Smith and Le (2017) empirically investigated how the optimal learning rate changes based on momentum and training set size. Large-batch training for faster convergence over multiple epochs has also been explored, with You et al. (2017) and Goyal et al. (2018) investigating methods to accelerate training by increasing batch sizes. Theoretical work has further sought to characterize CBS by analyzing SGD behavior in least-squares linear regression, especially in over-parameterized settings (Jain et al., 2018; Ma et al., 2018). Our work advances the optimization literature by formalizing CBS and quantifying its growth w.r.t. data size and emphasizing the importance of common hyper-parameter choices. It also provides strategies for studying large-scale pre-training beyond fixed training durations.
@@ 6 Concluding Remarks
In conclusion, this study provides an extensive examination of the scaling laws for critical batch size in large-scale autoregressive language model pre-training. By systematically analyzing the relationship between model size, data size, and CBS, we found that while CBS slightly increases with data size, it remains relatively invariant to model size. Our findings contribute insights into optimizing data parallelism and computational efficiency in pre-training setups, particularly highlighting the role of hyper-parameters such as learning rate scheduling and optimizer settings. These insights offer practical strategies for scaling models while maintaining efficiency, which is critical in resource-constrained scenarios. We leave for future work the investigation of scenarios involving more heterogeneous data sources than C4, as well as cases where the ratio of model size to data size substantially exceeds the Chinchilla-optimal configuration.
@@ Acknowledgements
We thank Dhruv Madeka. We thank Qirong Ho for discussions on distributed training; Jeremy Berstein for discussions on optimization; Dirk Groeneveld and Luca Soldaini for discussions on OLMo implementations. HZ is supported by an Eric and Susan Dunn Graduate Fellowship. SK acknowledges the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence; support from the Office of Naval Research under award N00014-22-1-2377, and the National Science Foundation Grant under award #IIS 2229881.
@@ Appendix A Training Dynamics
A simple strategy for setting warmup steps. To further prove that the critical batch size actually exists and the flattening of large batch sizes is not an artifact of not training well with proper hyper-paragrams, we take into account the warmup fraction in training as well:
@@
Scheduler comparison for other batch sizes. In Figure 10, we include more results on comparing different schedulers for batch sizes smaller than 8192 that are reported in the main text Figure 1. Overall, our Constant+EWA performs competitively with cosine scheduling and outperforms WSD scheduling, especially for large batch size regimes. Note that we sweep over the decay steps as 0.1,0.2,0.3×0.1,0.2,0.3$\times$0.1 , 0.2 , 0.3 × total training steps for WSD scheduling. Cosine scheduling also serves as a well-tuned baseline as we conduct sweeps over various maximum optimization steps to identify the optimal value, and then rerun the training using this step count. This approach ensures that the model reaches the target loss near the end of training, optimizing the performance of learning rate decay. Under small batch sizes, the schedule-free optimizer (Defazio et al., 2024) is a competitive baseline but it is significantly worse for batch sizes larger than 1024.
@@
Longer training requires higher EWA decay rate $\tau$. Throughout the paper, we adopt a learning rate of $0.00316$ across most experiments, but it is unclear whether that would be sub-optimal, especially since training with longer duration may require a lower learning rate as suggested in (DeepSeek-AI et al., 2024). Therefore, we justify our design decision on tuning EWA decay rate for simulating learning rate decay on different training durations by conducting the following experiments on a series of 151M models with (a) batch size 256, 0.5$\times$ Chinchilla tokens; (b) batch size 256, 20$\times$ Chinchilla tokens; (c) batch size 2048, 20$\times$ Chinchilla tokens, all with learning rate swept over $\{0.00316,0.00158,0.01264,0.00632,0.00075\}$ and EWA decay rate $\tau$ over $\{0.99,0.9968,0.999,0.99968,0.9999\}$. We set the number of warmup steps to be 0.25 of the total steps for (a) and 0.05 for (b) and (c). The results in Figure 11 denote the validation loss at the end of training, which consistently show that within each group of experiments, a learning rate of $0.00316$ we use throughout the paper is consistently the best. Moreover, when enlarging the training data size from 0.5$\times$ Chinchilla tokens to 20$\times$, the optimal EWA decay rate value $\tau$ would increase as well. This is also justified in the results presented in Figure 9 and Table 4 which indicate that longer training durations may benefit from a higher EWA decay rate to improve optimization performance.
Figure 11: Impact of learning rate and EWA decay rate across various training durations. Validation loss at the end of training for each hyper-parameter combination. $N$ denotes the model size. The best loss is marked by a
@@ Appendix D Additional Details on Experiments
Evaluation data size and frequency. As we need to frequently evaluate the model on a holdout evaluation set, understanding how large evaluation set size can lead to a stable and reliable measure yet still pertain efficiency of each run can be important. Therefore, we test out the evaluation variance on 327680 tokens is 2.17e-4, the variance when evaluating on 1638400 tokens is 4.53e-5, evaluation variance on 3276800 tokens is 7.65e-6. Therefore, we set the number of evaluation batches to be 100 by default. Note that the total number of training steps vary dramatically across batch sizes, so we adopt a hybrid evaluation protocol: the model is evaluated every $2^{i}$ ($i\in\mathbb{Z}$), every 1k steps and every $0.7n,0.75n,0.8n,\dots,n$ steps the last 30% of the total steps $n$. In this way, evaluation is conducted more frequently during the end of the training, thereby allowing a more accurate accounting of the total number of training steps required to get to a target evaluation loss.

### v2

@@ 4.1 Fixed Data Size and Scaling Up Model Size
Various previous works have established infinite width limits of neural networks (Yang and Hu, 2021; Bordelon and Pehlevan, 2022). For initializations and architectures that obey these limits, we can theoretically claim, that for a fixed training duration and batch size, the performance of the neural networks asymptotes with increasing width. The formal statement is provided below:
@@
For SGD with a given batch size $B$ (or for gradient descent, i.e., $B\rightarrow\infty$), training iterations $t$, an error tolerance $\epsilon>0$, fixed learning rate schedule and data ordering, for any network and initialization satisfying Master Theorem (Theorem G.4) in Yang and Hu (2021), there exists a width $w$ such that for any two networks $M_{1},M_{2}$ having widths $w_{1},w_{2}>w$, $|\mathcal{R}(M_{1},t)-\mathcal{R}(M_{2},t)|\leq\epsilon$, where $\mathcal{R}(M,t)$ denotes the loss of network $M$ at time $t$.
@@ 4.2 Fixed Model Size and Scaling up Data Size
We now turn to studying the impact of data size in mini-batch SGD for a well-specified Gaussian linear regression problem. Let $(\mathbf{x},y)$ be a pair of covariates and responses from a population distribution. Let the population risk and the population distribution be
@@
where $\gamma>0$ is a constant learning rate, $B$ is the batch size, $n:=D/B$ is the number of steps, $\mathbf{w}_{0}=0$ is the initialization (without loss of generality), and the output is the average of the iterates, $\bar{\mathbf{w}}:=\frac{1}{n}\sum_{t=0}^{n-1}\mathbf{w}_{t}$. Then the following theorem provides a tight bound on the excess risk achieved by the average of the mini-batch SGD iterates.
We write $f(D)\lesssim g(D)$ if there is a positive constant $c$ such that $f(D)\leq cg(D)$ for every $D\geq 1$. We write $f(D)\eqsim g(D)$ if $f(D)\lesssim g(D)\lesssim f(D)$. The proofs are all deferred to Appendix G.
@@
Let $(\lambda_{i})_{i>0}$ be the eigenvalues of $\mathbf{H}$ in nonincreasing order. Assume that $\|\mathbf{w}_{0}-\mathbf{w}^{*}\|_{\mathbf{H}}^{2}\lesssim\sigma^{2}$. Then for every $\gamma\lesssim\min\{B/\mathrm{tr}(\mathbf{H}),1/\|\mathbf{H}\|_{2}\}$, we have
where $k^{*}:=\max\{k:\lambda_{k}\geq B/(D\gamma)\}$ and the expectation is over the randomness of $\bar{\mathbf{w}}$.
The proof of Theorem 3 is motivated by Zou et al. (2023). We focus on well-specified Gaussian data distribution for simplicity, but this can be relaxed to misspecified cases under fourth-moment conditions following the results in Zou et al. (2023). Theorem 3 suggests that for a fixed data size $D$, the excess risk depends on the batch size $B$ and the learning rate $\gamma$ only through their ratio $\gamma/B$. Moreover, a large $\gamma/B$ tends to decrease the bias error (the terms depending on $\mathbf{w}^{*}$) but increase the variance error (the terms depending on $\sigma^{2}$), and vice versa. This observation is exploited in the following corollary, where we compute the CBS that minimizes the sequential running time without sacrificing the rate of the attained excess risk.
Corollary 2.
Under the settings of Theorem 3, additionally assume $\sigma^{2}\eqsim 1$ and the following capacity and source conditions:
@@
  1. 1.
  2. 2.
    When $b>a$, the optimal hyper-parameters are $\gamma^{*}$ and $B^{*}$ such that
    Therefore, the CBS is $B^{*}\eqsim D^{1-a/\min\{b,2a+1\}}$, which (along with $\gamma^{*}\eqsim 1$) allows mini-batch SGD output $\bar{\mathbf{w}}$ to attain the optimal rate of the expected excess risk (as data size $D$ grows) with the smallest number of steps $n$.
The capacity and source conditions are from the nonparametric linear regression literature (Caponnetto and De Vito, 2007) and are recently used to study scaling laws theory (Bordelon et al., 2024c; Lin et al., 2024; Paquette et al., 2024). According to Corollary 2, when $b\leq a$, the bias error tends to dominate the variance error, in this case, the CBS is $B^{*}=1$ to allow a maximum number of optimization steps. when $b<a$, the variance error tends to dominate the bias error, and the optimal choices of batch size and learning rate balance these two errors. While one can use $B^{*}=1$ and a small $\gamma^{*}$ to achieve the best excess risk rate, this leads to a suboptimal sequential runtime ($n=D/B$). In this case, the CBS is $B^{*}\eqsim D^{1-a/\min\{b,2a+1\}}$, which achieves the optimal excess risk rate while minimizing the sequential runtime.
@@ 5 Related Work
Scaling laws describe the parametric relationships among key factors involved in training neural networks: model size $N$, dataset size $D$, training cost $C$, and final training loss $\mathcal{R}$. These laws enable the prediction of training loss $\mathcal{R}$ based on available resources, making it possible to optimize resource allocation for efficient model training. For example, Hestness et al. (2017) found $\mathcal{R}\propto D^{-\alpha}$, with $\alpha\in[0.07,0.35]$. Of the factors they varied, only tasks can change the exponent $\alpha$. Changing the architecture optimizers, regularizers, and loss functions, would only change the proportionality factor, not the exponent; Henighan et al. (2020) studied statistical relations between $N,D,C,\mathcal{R}$, over a wide range of values and found similar scaling laws, over the range of $N\in\left[10^{3},10^{9}\right],C\in\left[10^{12},10^{21}\right]$, and over multiple modalities (text, video, image, text to image, etc.). (Kaplan et al., 2020) states that $N$ should be scaled faster than $D$. However, Chinchilla scaling (Hoffmann et al., 2022b) found that models are under-trained, and then suggests that when given an increased budget (in FLOPs), to achieve compute-optimal, model size $N$ and data size $D$ should scale in approximately equal proportions. Recent efforts (Pearce and Song, 2024; Besiroglu et al., 2024; Porian et al., 2024) have been made in reproducing the scaling laws from (Hoffmann et al., 2022b) and the (Kaplan et al., 2020). Different from our focus on measuring the efficiency notion of CBS, most of them focus on deriving optimal hyper-parameters (Bi et al., 2024; Porian et al., 2024) including learning rate and batch size from small-scale training given a fixed compute budget FLOPs $\approx 6ND$ without decoupling the effects of model size and data size.
Optimization and critical batch size.
@@
Previous studies have shown that increasing batch sizes can be offset by a proportional adjustment to the learning rate in small-scale regimes (McCandlish et al., 2018; Zhang et al., 2019; Kaplan et al., 2020). McCandlish et al. (2018) introduce the gradient noise scale, a measure that captures the variation in gradients across different training examples, which helps predict the critical batch size (CBS). Their findings also suggest that small-batch training is more compute-efficient, while large-batch training requires fewer optimizer steps. Momentum-based methods extend scaling to larger batch sizes but converge to the performance of standard SGD at smaller batch sizes (Shallue et al., 2019). Additionally, Zhang et al. (2019) analyze the impact of curvature on CBS using a noisy quadratic model, demonstrating that preconditioning techniques can increase the CBS. Golmant et al. (2018) show that the size of the dataset plays a smaller role in determining training efficiency compared to factors like model architecture and data complexity. In contrast, Hilton et al. (2022) examine how performance can be maintained at smaller batch sizes. Meanwhile, Smith et al. (2017); Smith and Le (2017) empirically investigated how the optimal learning rate changes based on momentum and training set size. Theoretical work has further sought to characterize CBS by analyzing SGD behavior in least-squares linear regression, especially in over-parameterized settings (Jain et al., 2018; Ma et al., 2018). Filatov et al. (2024) concurrently find that optimal batch size and CBS scale with data size. However, they do not explore how CBS scales with model size for models beyond 354M parameters, nor do they provide theoretical justifications or address the challenge of selecting optimal runs across a broad range of hyperparameters. Our work advances the optimization literature by formalizing CBS and quantifying its growth w.r.t. data size and emphasizing the importance of common hyper-parameter choices. It also provides strategies for studying large-scale pre-training beyond fixed training durations.
@@ 6 Concluding Remarks
In conclusion, this study provides an extensive examination of the scaling laws for critical batch size in large-scale autoregressive language model pre-training. By systematically analyzing the relationship between model size, data size, and CBS, we found that while CBS increases with data size, it remains relatively invariant to model size. This finding suggests training on more data may enable greater data parallelism in pre-training. We further emphasize the role of key hyperparameters and exponential weight averaging, which can match the performance cosine scheduling without requiring fixed training durations. These insights offer practical strategies for scaling models while maintaining efficiency, which is critical in resource-constrained scenarios.
@@ Acknowledgements
We thank Dhruv Madeka. We thank Qirong Ho for discussions on distributed training; Jeremy Berstein for discussions on optimization; Dirk Groeneveld and Luca Soldaini for discussions on OLMo implementations. HZ is supported by an Eric and Susan Dunn Graduate Fellowship. SK and DM acknowledge the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence; SK acknowledges the support from the Office of Naval Research under award N00014-22-1-2377, and the National Science Foundation Grant under award #IIS 2229881. NV and DM are supported by a Simons Investigator Fellowship, NSF grant DMS-2134157, DARPA grant W911NF2010021, and DOE grant DE-SC0022199.
@@ Appendix A Training Dynamics
A simple strategy for setting warmup steps. To further prove that the critical batch size actually exists and the saturation of large batch sizes is not an artifact of not training well with proper hyper-paragrams, we take into account the warmup fraction in training as well:
@@
Scheduler comparison for other batch sizes. In Figure 9, we include more comparisons on different schedulers that are reported in the main text (Figure 1). Overall, our Constant+EWA performs competitively with cosine scheduling and outperforms WSD scheduling, especially for large batch size regimes. Note that we sweep over the decay steps as 0.1,0.2,0.3×0.1,0.2,0.3$\times$0.1 , 0.2 , 0.3 × total training steps for WSD scheduling. We tune cosine scheduling by conducting sweeps over various maximum optimization steps to identify the optimal value, and then rerun the training using this step count. This approach ensures that the model reaches the target loss near the end of training, optimizing the performance of learning rate decay. Under small batch sizes, the schedule-free optimizer (Defazio et al., 2024) is a competitive baseline but it is significantly worse for batch sizes larger than 1024.
@@
Longer training requires higher EWA decay rate $\tau$. Throughout the paper, we adopt a learning rate of $0.00316$ across most experiments, but it is unclear whether that would be sub-optimal, especially since training with longer duration may require a lower learning rate as suggested in (DeepSeek-AI et al., 2024). Therefore, we justify our design decision on tuning EWA decay rate for simulating learning rate decay on different training durations by conducting the following experiments on a series of 151M models with (a) batch size 256, 0.5$\times$ Chinchilla tokens; (b) batch size 256, 20$\times$ Chinchilla tokens; (c) batch size 2048, 20$\times$ Chinchilla tokens, all with learning rate swept over $\{0.00316,0.00158,0.01264,0.00632,0.00075\}$ and EWA decay rate $\tau$ over $\{0.99,0.9968,0.999,0.99968,0.9999\}$. We set the number of warmup steps to be 0.25 of the total steps for (a) and 0.05 for (b) and (c). The results in Figure 10 denote the validation loss at the end of the training, which consistently show that within each group of experiments, a learning rate of $0.00316$ we use throughout the paper is consistently the best. Moreover, when enlarging the training data size from 0.5$\times$ Chinchilla tokens to 20$\times$, the optimal EWA decay rate value $\tau$ would increase as well. This is also justified in the results presented in Figure 8 and Table 4 which indicate that longer training durations may benefit from a higher EWA decay rate to improve optimization performance.
Figure 10: Impact of learning rate and EWA decay rate across various training durations. We report the validation loss at the end of training for each hyper-parameter combination. $N$ denotes the model size. The best loss is marked by a
@@ Appendix C Results Including Small Batch Sizes
Furthermore, recall that we set $B_{\text{opt}}=2^{8}$ when selecting the target loss. Figure 12 and Figure 13 confirm that $2^{8}$ falls within the linear scaling regime, which justifies our design choice.
@@ Appendix D Additional Details on Experiments
Evaluation data size and frequency. As we need to frequently evaluate the model on a holdout evaluation set, understanding how large evaluation set size can lead to a stable and reliable measure yet still pertain efficiency of each run can be important. We evaluated the variance across different token counts: 2.17e-4 for 327,680 tokens, 4.53e-5 for 1,638,400 tokens, and 7.65e-6 for 3,276,800 tokens. Based on these results, we have set the default number of evaluation batches to 100.
It is important to note that the total number of training steps varies significantly across different batch sizes. To address this, we implement a hybrid evaluation protocol: the model is evaluated at intervals of $2^{i}$ (where $i\in\mathbb{Z}$ ), every 1,000 steps, and at $0.7n,0.75n,0.8n,\ldots$, up to $n$ during the last $30\%$ of the total steps $n$. This approach ensures more frequent evaluations toward the end of training, allowing for a more accurate assessment of the total training steps needed to achieve a target evaluation loss.